OpenStudy (anonymous):

How do you find y' for the given equation? y = x-2/x

OpenStudy (anonymous):

Quotient rule where u = x-2 and v = x \[\frac{ du }{dv } \frac{ u }{ v } = \frac{ u'v-v'u }{ v^2}\]

OpenStudy (precal):